# Continuous in time bubbling and Soliton Resolution for Non-negative Solutions of the Energy-Critical Heat Flow

### v1

@@ 1.1. Problem Setting
which appears naturally in the study of extremizers of the Sobolev inequality and, more generally, is connected to the Yamabe problem on the sphere via stereographic projection. The local well-posedness of (1.1) in $\dot{H}^{1}$-norm is classical and was initiated by Weissler in [W1, W2], with further contributions by Giga [giga-3], Ni–Sacks [ni-sacks], and Brezis-Cazenave [brezis-cazenave]. Observe that the solutions of (1.1) are invariant under translations and parabolic scaling,
@@
Since the nonlinear energy is invariant under these symmetries, i.e., $E(u)=E\left(u_{\lambda}\right)$, the equation (1.1) is energy-critical. Testing (1.1) against $\partial_{t}u$ and integrating by parts we observe the formal energy identity
@@
1.2. Statement of the Main result
To state our main Theorem 1.6, we first define a notion of scale and center of a non-zero stationary solution. Let $S=S(d)>0$ be the sharp constant in the Sobolev inequality on $\mathbb{R}^{d}$,
@@ 1.2. Statement of the Main result
By the variational characterization of the Sobolev inequality, the best constant (or equality) is attained by a positive stationary solution $W^{\prime}$, i.e.,
$\begin{split}\int_{\mathbb{R}^{d}}|W^{\prime}|^{p+1}\mathrm{d}x=S^{p+1}\left(\int_{\mathbb{R}^{d}}|\nabla W^{\prime}|^{2}\mathrm{d}x\right)^{(p+1)/2}=\int_{\mathbb{R}^{d}}|\nabla W^{\prime}|^{2}\mathrm{d}x\end{split}$ (1.8)
@@
$\begin{split}\int_{\mathbb{R}^{d}}|\nabla W^{\prime}|^{2}\mathrm{d}x=S^{-d}.\end{split}$ (1.9)
@@
In particular, we deduce that any non-zero stationary solution $W\in\dot{H}^{1}$ satisfies
$\|\nabla W\|^{2}_{L^{2}}\geq S^{-d}.$ (1.11)
Denote $\bar{E}_{*}:=S^{-d}$ as the minimal energy of any non-zero stationary solution of (1.5), and in general, define $\bar{E}(u;\Omega):=\|\nabla u\|_{L^{2}(\Omega)}^{2}$ for any $\Omega\subset\mathbb{R}^{d}$ and $u\in\dot{H}^{1}$ with the convention that $\bar{E}(u):=\bar{E}(u;\mathbb{R}^{d}).$ Thus, given any non-zero stationary solution $W:\mathbb{R}^{d}\to\mathbb{R}$, we define its scale and center as follows:
@@
where $W_{j}:\mathbb{R}^{d}\to\mathbb{R}$ are smooth non-zero stationary solutions. By convention if $K=0$ then $\mathbf{W}\equiv 0$. To emphasize the dependence of $\mathbf{W}$ on the collection $\{W_{j}\}_{j=1}^{K}$, we will occasionally write $\mathbf{W}=\mathbf{W}(\vec{W})$, where $\vec{W}=(W_{1},\ldots,W_{K}).$
@@
Given, $y\in\mathbb{R}^{d}$, $\rho\in(0,\infty)$, $u:[0,T_{+})\times B(y,\rho)\to\mathbb{R}$, where $T_{+}>0$ and $\gamma_{0}\in(0,\bar{E}_{*}/2)$, define
@@
where the infimum above is taken over all possible $K$-multi bubble configurations for any non-negative integer $K$, over all parameters $\vec{\nu}\in(0,\infty)^{K+1}$ and $\vec{\xi}\in(0,\infty)^{K+1}$ as in Definition 1.4. Since we will fix $\gamma_{0}$ later, we drop the subscript involving $\gamma_{0}$ in subsequent expressions.
With these definitions in hand, we state the main theorem in this paper.
Theorem 1.6 (Continuous Bubbling for NLH).
Let $u(t)$ be a solution of (1.1) with initial data $u_{0}\in\dot{H}^{1}$. Let $T_{+}=T_{+}(u_{0})\in(0,\infty]$ denote its maximal time of existence and assume that $u(t)$ has finite energy, i.e., $\sup_{t\in[0,T_{+})}\bar{E}(u(t))<\infty.$ Then the following hold:
$\mathrm{(i)}$ If $T_{+}<\infty$, then there exists a finite energy map $u^{*}:\mathbb{R}^{d}\to\mathbb{R}$, an integer $K\geq 1$, and points $\{x^{i}\}_{i=1}^{K}\subset\mathbb{R}^{d}$ such that the following holds: let $t_{n}\to T_{+}$ be any time sequence. After passing to a subsequence (still denoted by $t_{n}$) we can associate to each $i\in\{1,\dots,K\}$ an integer $J_{i}$, sequences $a_{j,n}^{i}\in\mathbb{R}^{d}$ and $\lambda_{j,n}^{i}\in(0,\infty)$ for each $j\in\{1,\dots,J_{i}\}$, with $a_{j,n}^{i}\to x^{i}$, $\frac{\lambda_{j,n}^{i}}{\sqrt{T_{+}-t_{n}}}\to 0$ as $n\to\infty$, and non-zero bubbles $W_{1}^{i},\dots,W_{J_{i}}^{i}$ such that
and
and non-zero bubbles $W_{1},\dots,W_{K}$, so that
and
Note that the bubbles obtained in the above decomposition may depend on the sequence of times, which is similar to an issue encountered in [lawrie-harmonic-map-nonradial]. However, we can resolve this issue for (1.1) with a very reasonable assumption, yielding the full Soliton Resolution Conjecture in the case of finite energy non-negative initial data. See also the recent work [ishiwata2018potential] where a similar result is proved by different methods in the case when $T_{+}=\infty.$
Corollary 1.7.
Let $u(t)$ be a solution of (1.1) with non-negative initial data $u_{0}\geq 0$ and $u_{0}\in\dot{H}^{1}$. Let $T_{+}=T_{+}(u_{0})\in(0,\infty]$ denote its maximal time of existence and assume that $u(t)$ has finite energy, i.e., $\sup_{t\in[0,T_{+})}\bar{E}(u(t))<\infty.$ Then the maps obtained in the decompositions (1.18) and (1.21) are unique and independent of the sequence of times.
Theorem 1.6 is a consequence of the following localized bubbling result, in which we denote the ball centered at $x\in\mathbb{R}^{d}$ with radius $r>0$ as $B(x,r):=\{y\in\mathbb{R}^{d}:|x-y|<r\}.$
@@ 1.3. Background and Motivation
A fundamental problem in the analysis of nonlinear partial differential equations (PDEs) is describing the long-time behavior of their solutions. The Soliton Resolution Conjecture asserts that any finite-energy solution to a dispersive PDE asymptotically decomposes into a sum of decoupled solitons that are stationary solutions of the underlying equation, a radiation term that behaves like a solution to the linear flow, and an error term that vanishes in the natural energy norm. This conjecture arose from the numerical experiments of Fermi–Pasta–Ulam–Tsingou [pasta] and Zabusky–Kruskal [zabusky], which provided evidence that it holds for the Korteweg-de Vries (KdV) equation. Since then, the problem has been extensively studied for the KdV equation as well as for several other integrable models.
Beyond integrable systems, analogues of the Soliton Resolution Conjecture have emerged across various areas of mathematics. In general relativity, the Final State Conjecture (cf. [Klainerman]) predicts that generic solutions to Einstein’s field equations asymptotically approach a finite number of stationary solutions or Kerr black holes moving apart from each other. In geometric analysis, Soliton Resolution arises naturally in the study of gradient flows associated with conformally invariant variational problems. For example, pioneering works of Struwe [struwe, struwe1994yang], Qing [qing], Qing–Tian [qing-tian], and Hong–Tian [hong-tian] have established bubbling or Soliton Resolution along a well-chosen sequence of times for the harmonic map and Yang–Mills heat flows.
Motivated by these parabolic works, in this paper, we study the energy-critical nonlinear heat flow in dimension $d\geq 3.$ Our main result, Theorem 1.6, establishes a continuous-in-time bubble-tree decomposition for all finite-energy solutions of (1.1). More precisely, any solution with uniformly bounded $\dot{H}^{1}$-norm decomposes into a sum of solitons that may vary along different time sequences, a radiation term that is asymptotically trivial or captured by a weak limit in $\dot{H}^{1}$, and an error term that vanishes in the energy space. Moreover, when the initial data is non-negative, Corollary 1.7 shows that Theorem 1.6 implies the Soliton Resolution Conjecture, since positive solitons have been classified and are unique up to the symmetries of the equation due to [obata1972conformal, CGS]. Therefore, Theorem 1.6 extends Struwe’s classical compactness result [struwe-global], which establishes a similar decomposition only along a well-chosen sequence of times, while Corollary 1.7 provides the first instance of Soliton Resolution for a non-integrable PDE, beyond radial symmetry, and without restrictions on the size of the initial data.
@@
For non-integrable equations with radial symmetry, where the solitons do not move in space, the conjecture has been settled for the nonlinear wave equation [Duyckaerts-wave:d=3, Duyckaerts-wave-2, Duyckaerts-wave-5, Duyckaerts-wave-6, jia-kenig, collot2022soliton, lawrie-wave, lawrie-wavemaps], damped Klein–Gordon equation [burq2017long, gu2023soliton], equivariant self-dual Chern–Simons–Schrödinger equation [oh-chern-simons], equivariant harmonic map heat flow [lawrie-harmonic-map], and energy-critical nonlinear heat flow [nlh-radial].
For non-integrable equations, without radial symmetry, Soliton Resolution is known in one dimension for the damped Klein-Gordon equation [cote2021long], in the neighborhood of a few solitons for the energy-critical nonlinear heat flow and the damped Klein-Gordon equation [collot, ishizuka2023global, ishizuka2025long], continuously in time for the harmonic map heat flow [lawrie-harmonic-map-nonradial] or along a sequence of times in dimensions $3\leq d\leq 5$ for the energy-critical nonlinear wave equation [jia-dkm].
In contrast, establishing our main results requires working in any dimension $d\geq 3$, where solitons exhibit only weak decay and no longer enjoy radial symmetry, allowing them to translate in space and potentially behave pathologically (cf. [ding1986conformally, del2011large, del2013torus]). Moreover, we impose no restrictions on the size of the initial data, which implies that the nonlinear energy is, in general, non-coercive, unlike the setting of [kenig-merle, roxanas]. We overcome these difficulties by introducing new ideas that are robust and adaptable to other nonlinear parabolic flows, as we will show in forthcoming works. In particular, our modified notion of collision intervals, introduced in Section 3, can be used to generalize the results of [lawrie-harmonic-map-nonradial] to higher-dimensional target manifolds.
@@ 1.4. Proof Sketch
The proofs of the main Theorems 1.6 and 1.8 build on the framework of [lawrie-harmonic-map-nonradial], but require addressing new difficulties that arise in the context of the energy-critical nonlinear heat flow. This includes:
@@
Therefore, to show that no energy lives outside the ball $B(y,\sqrt{t})$ for any $y\in\mathbb{R}^{d}$, it suffices to show that the localized nonlinear energy and the $L^{p+1}$-norm vanish in this region. By localizing (1.4), it is not difficult to show that the first quantity vanishes. On the other hand, the vanishing of $L^{p+1}$ norm is quite involved and, in particular, relies on a recent result of Ishiwata [ishiwata2018potential], see Lemma 2.16.
Now let $t_{n}\to T_{+}$ be any sequence of times. From Theorem 1.8, we see that $u(t_{n})$ approaches $K$ multi-bubble configuration on either $B(y,\sqrt{T_{+}-t})$ when $T_{+}<\infty$ or $B(y,\sqrt{t})$ when $T_{+}=\infty.$ In particular, the $K$ multi-bubble configurations depend on $n$. To obtain a finite number of bubbles (independent of $n$) as in Theorem 1.6 that are asymptotically orthogonal in the sense of (1.17) and (1.19), we apply the Compactness Theorem 2.18 to each bubble obtained in the sequence of multi-bubble configurations arising from Theorem 1.8 and build a new bubble tree configuration by selecting bubbles such that (1.17) and (1.19) are satisfied. The resulting multi-bubble configuration then satisfies all the requirements of Theorem 1.6, thus completing the proof.

### v2

@@ 1.1. Problem Setting
which arises naturally in the study of extremizers of the Sobolev inequality and, more generally, is related to the Yamabe problem on the sphere via stereographic projection. The local well-posedness of (1.1) in $\dot{H}^{1}$-norm is classical and was initiated by Weissler in [W1, W2], with further contributions by Giga [giga-3], Ni–Sacks [ni-sacks], and Brezis-Cazenave [brezis-cazenave]. Observe that the solutions of (1.1) are invariant under translations and parabolic scaling,
@@
Since the nonlinear energy is invariant under these symmetries, i.e., $E(u)=E\left(u_{\lambda}\right)$, the equation (1.1) is energy-critical. Testing (1.1) against $\partial_{t}u$ and integrating by parts, we observe the formal energy identity
@@
1.2. Statement of the Main Results
We now state our first main result.
Theorem 1.1 (Continuous Bubbling for NLH).
Let $u(t)$ be a solution of (1.1) with initial data $u_{0}\in\dot{H}^{1}$. Let $T_{+}=T_{+}(u_{0})\in(0,\infty]$ denote its maximal time of existence and assume that $u(t)$ has finite energy, i.e., $\sup_{t\in[0,T_{+})}\|\nabla u(t)\|_{L^{2}}<\infty.$ Then the following hold:
$\mathrm{(i)}$ If $T_{+}<\infty$, then there exist a finite energy map $u^{*}\in\dot{H}^{1}$, an integer $K\geq 1$, and points $\{x^{i}\}_{i=1}^{K}\subset\mathbb{R}^{d}$ such that the following holds. Let $t_{n}\to T_{+}$ be any time sequence. After passing to a subsequence (still denoted by $t_{n}$) we can associate to each $i\in\{1,\dots,K\}$ an integer $J_{i}$, sequences $a_{j,n}^{i}\in\mathbb{R}^{d}$ and $\lambda_{j,n}^{i}\in(0,\infty)$ for each $j\in\{1,\dots,J_{i}\}$, with $a_{j,n}^{i}\to x^{i}$, $\frac{\lambda_{j,n}^{i}}{\sqrt{T_{+}-t_{n}}}\to 0$ as $n\to\infty$, and non-zero bubbles $W_{1}^{i},\dots,W_{J_{i}}^{i}$ such that
and
and non-zero bubbles $W_{1},\dots,W_{K}$, so that
and
Remark 1.2.
Note that since our arguments do not rely on modulation techniques, we do not obtain a continuous family of scaling and translation parameters as done, for instance, in the radial case (cf. Theorem 1.1 in [nlh-radial]).
Observe that the bubbles obtained in the above decomposition may depend on the sequence of times, which remains a challenging issue to be understood even in the case of the harmonic map heat flow (cf. Remark 1.8 in [lawrie-harmonic-map-nonradial]). However, we can resolve this issue for (1.1) with a very reasonable assumption, yielding the full Soliton Resolution Conjecture in the case of finite energy non-negative initial data. Before stating the result, we recall that by the work of [CGS], any positive solution with finite $\dot{H}^{1}$-norm of the elliptic problem (1.5) is given by the so-called Aubin–Talenti bubbles, defined as maps $\widetilde{W}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ such that
$\widetilde{W}(x):=\left(1+\frac{|x|^{2}}{d(d-2)}\right)^{-\frac{d-2}{2}}$ (1.11)
up to scaling, and translation. This elliptic rigidity is enough to enforce uniqueness of the limiting bubbles in Theorem 1.1. Consequently, we obtain
Corollary 1.3.
Let $u(t)$ be a solution to (1.1) with non-negative initial data $u_{0}\geq 0$ and $u_{0}\in\dot{H}^{1}$. Let $T_{+}=T_{+}(u_{0})\in(0,\infty]$ denote its maximal time of existence and assume that $u(t)$ has finite energy, i.e., $\sup_{t\in[0,T_{+})}\|\nabla u(t)\|_{L^{2}}<\infty.$ Then the bubbles obtained in the decompositions (1.7) and (1.10) are unique and are given by the Aubin–Talenti bubbles defined in (1.11).
Theorem 1.1 follows from a localized bubbling result that requires a few definitions. We first define a notion of scale and center of a non-zero stationary solution. Let $S=S(d)>0$ be the sharp constant in the Sobolev inequality on $\mathbb{R}^{d}$,
@@ 1.2. Statement of the Main Results
By the variational characterization of the Sobolev inequality, the best constant (or equality) is attained by a positive stationary solution $\widetilde{W}$, i.e.,
$\begin{split}\int_{\mathbb{R}^{d}}|\widetilde{W}|^{p+1}\mathrm{d}x=S^{p+1}\left(\int_{\mathbb{R}^{d}}|\nabla\widetilde{W}|^{2}\mathrm{d}x\right)^{(p+1)/2}=\int_{\mathbb{R}^{d}}|\nabla\widetilde{W}|^{2}\mathrm{d}x\end{split}$ (1.14)
@@
$\begin{split}\int_{\mathbb{R}^{d}}|\nabla\widetilde{W}|^{2}\mathrm{d}x=S^{-d}.\end{split}$ (1.15)
@@
In particular, we deduce that any non-zero stationary solution $W\in\dot{H}^{1}$ satisfies $\|\nabla W\|^{2}_{L^{2}}\geq S^{-d}.$ Denote $\bar{E}_{*}:=S^{-d}$ as the minimal energy of any non-zero stationary solution of (1.5), and in general, define $\bar{E}(u;\Omega):=\|\nabla u\|_{L^{2}(\Omega)}^{2}$ for any $\Omega\subset\mathbb{R}^{d}$ and $u\in\dot{H}^{1}$ with the convention that $\bar{E}(u):=\bar{E}(u;\mathbb{R}^{d}).$ Let $B(x,r)\subset\mathbb{R}^{d}$ denote the ball centered at $x\in\mathbb{R}^{d}$ with radius $r>0.$ Given any non-zero stationary solution $W:\mathbb{R}^{d}\to\mathbb{R}$, we define its scale and center as follows:
@@
where $W_{j}:\mathbb{R}^{d}\to\mathbb{R}$ are smooth non-zero stationary solutions. By convention, if $K=0$ then $\mathbf{W}\equiv 0$. To emphasize the dependence of $\mathbf{W}$ on the collection $\{W_{j}\}_{j=1}^{K}$, we will occasionally write $\mathbf{W}=\mathbf{W}(\vec{W})$, where $\vec{W}=(W_{1},\ldots,W_{K}).$
@@
Given, $y\in\mathbb{R}^{d}$, $\rho\in(0,\infty)$, $u:[0,T_{+})\times\mathbb{R}^{d}\to\mathbb{R}$, where $T_{+}>0$ and $\gamma_{0}\in(0,\bar{E}_{*}/2)$, define
@@
where the infimum above is taken over all possible $K$ multi-bubble configurations for any non-negative integer $K$, over all parameters $\vec{\nu}\in(0,\infty)^{K+1}$ and $\vec{\xi}\in(0,\infty)^{K+1}$ as in Definition 1.7. Since we will fix $\gamma_{0}$ later, we drop the subscript involving $\gamma_{0}$ in subsequent expressions.
We are now ready to state our second main result, which implies Theorem 1.1.
@@ 1.3. Background and Motivation
A fundamental problem in the analysis of nonlinear partial differential equations (PDEs) is describing the long-time behavior of their solutions. The Soliton Resolution Conjecture asserts that any finite-energy solution to a nonlinear PDE asymptotically decomposes into a sum of decoupled solitons that are stationary solutions of the underlying equation, a radiation term that behaves like a solution to the linear flow, and an error term that vanishes in the natural energy norm. This conjecture arose from the numerical experiments of Fermi–Pasta–Ulam–Tsingou [pasta] and Zabusky–Kruskal [zabusky], which provided evidence that it holds for the Korteweg-de Vries (KdV) equation. Since then, the problem has been extensively studied for the KdV equation as well as for several other integrable models.
Beyond integrable systems, analogues of the Soliton Resolution Conjecture have emerged across various areas of mathematics. In general relativity, the Final State Conjecture (cf. [Klainerman]) predicts that generic solutions to Einstein’s field equations asymptotically approach a finite number of stationary solutions, namely Kerr black holes, moving apart from each other. In geometric analysis, Soliton Resolution arises naturally in the study of gradient flows associated with conformally invariant variational problems. For example, pioneering works of Struwe [struwe, struwe1994yang], Qing [qing], Qing–Tian [qing-tian], and Hong–Tian [hong-tian] have established bubbling or Soliton Resolution along a well-chosen sequence of times for the harmonic map and Yang–Mills heat flows.
Motivated by these parabolic works, we study the energy-critical nonlinear heat flow in dimension $d\geq 3.$ Our main result, Theorem 1.1, establishes a continuous-in-time bubble-tree decomposition for all $\dot{H}^{1}$ bounded solutions of (1.1). More precisely, any solution with uniformly bounded $\dot{H}^{1}$-norm decomposes into a sum of solitons that may vary along different time sequences, a radiation term that is asymptotically trivial or captured by a weak limit in $\dot{H}^{1}$, and an error term that vanishes in the energy space. Moreover, when the initial data is non-negative, Corollary 1.3 shows that Theorem 1.1 implies the Soliton Resolution Conjecture, since positive solitons have been classified and are unique up to the symmetries of the equation due to [obata1972conformal, CGS]. Therefore, Theorem 1.1 extends Struwe’s classical compactness result [struwe-global], which establishes a similar decomposition only along a well-chosen sequence of times. It also strengthens Theorem 1.4 in [ishiwata2018potential], which proves a similar decomposition for global-in-time solutions with an error term vanishing in the $L^{p+1}$-norm. On the other hand, Corollary 1.3 establishes Soliton Resolution for a non-integrable PDE beyond radial symmetry and without restrictions on the size of the initial data. A similar result has been established by Ishiwata [ishiwata2018potential] for global in time solutions, whereas our work covers both finite-time blowup and the global-in-time regime.
@@
For non-integrable equations with radial symmetry, where the solitons do not move in space, the conjecture has been settled for the nonlinear wave equation [Duyckaerts-wave:d=3, Duyckaerts-wave-2, Duyckaerts-wave-5, Duyckaerts-wave-6, jia-kenig, collot2022soliton, lawrie-wave, lawrie-wavemaps], damped Klein–Gordon equation [burq2017long, gu2023soliton], equivariant self-dual Chern–Simons–Schrödinger equation [oh-chern-simons], equivariant harmonic map heat flow [lawrie-harmonic-map], and energy-critical nonlinear heat flow [nlh-radial]. The last two works led to the first complete classification of the bubble tree term dynamics for the equivariant harmonic map heat flow and the radial energy-critical nonlinear heat flow by Kim and Merle in [kim2024classification].
For non-integrable equations without radial symmetry, Soliton Resolution is known in one dimension for the damped Klein-Gordon equation [cote2021long], in the neighborhood of a few solitons for the energy-critical nonlinear heat flow and the damped Klein-Gordon equation [collot, ishizuka2023global, ishizuka2025long], continuously in time for the harmonic map heat flow [lawrie-harmonic-map-nonradial] or along a sequence of times in dimensions $3\leq d\leq 5$ for the energy-critical nonlinear wave equation [jia-dkm].
Regarding (1.1) in the energy-critical setting, Wang–Wei [wang-wei] proved that in dimensions $d\geq 7$ and for nonnegative initial data $u_{0}\geq 0$, any finite-time blowup must be of Type I, thereby ruling out finite-time Type II blowup in higher dimensions. When $3\leq d\leq 6$, examples of Type II blowup have been constructed, for instance by del Pino–Musso–Wei–Zhang–Zhang [del2020type], Schweyer [schweyer2012type], Pino–Musso–Wei [del2019type], and Harada [harada2020type].
In contrast to the existing work, establishing our main results requires working in any dimension $d\geq 3$, where solitons exhibit only weak decay and no longer enjoy radial symmetry, allowing them to translate in space and potentially behave pathologically (cf. [ding1986conformally, del2011large, del2013torus]). Moreover, we impose no restrictions on the size of the initial data, which implies that the nonlinear energy is, in general, non-coercive, unlike the setting of [kenig-merle, roxanas]. We overcome these difficulties by introducing new ideas that are robust and adaptable to other nonlinear parabolic flows, as we will show in forthcoming works. In particular, our modified notion of collision intervals, introduced in Section 3, can be used to generalize the results of [lawrie-harmonic-map-nonradial] to higher-dimensional target manifolds.
@@ 1.4. Proof Sketch
The proofs of the main Theorems 1.1 and 1.9 build on the framework of [lawrie-harmonic-map-nonradial], but require addressing new difficulties that arise in the context of the energy-critical nonlinear heat flow. This includes:
@@
Therefore, to show that no energy lives outside the ball $B(y,\sqrt{t})$ for any $y\in\mathbb{R}^{d}$, it suffices to show that the localized nonlinear energy and the $L^{p+1}$-norm vanish in this region. By localizing (1.4), it is not difficult to show that the first quantity vanishes. On the other hand, the vanishing of $L^{p+1}$ norm is quite involved and, in particular, relies on a deep work by Ishiwata [ishiwata2018potential], see Lemma 2.16.
Now let $t_{n}\to T_{+}$ be any sequence of times. From Theorem 1.9, we see that $u(t_{n})$ approaches $K$ multi-bubble configuration on either $B(y,\sqrt{T_{+}-t_{n}})$ when $T_{+}<\infty$ or $B(y,\sqrt{t}_{n})$ when $T_{+}=\infty.$ In particular, the $K$ multi-bubble configurations depend on $n$. To obtain a finite number of bubbles (independent of $n$) as in Theorem 1.1 that are asymptotically orthogonal in the sense of (1.6) and (1.8), we apply the Compactness Theorem 2.18 to each bubble obtained in the sequence of multi-bubble configurations arising from Theorem 1.9 and build a new bubble tree configuration by selecting bubbles such that (1.6) and (1.8) are satisfied. The resulting multi-bubble configuration then satisfies all the requirements of Theorem 1.1, thus completing the proof.